# Mechanistic Interpretability for AI Safety A Review

### v1

@@ 9 Future Directions
Given the current limitations and challenges, several promising directions can be pursued to advance mechanistic interpretability, emphasizing conceptual clarity, establishing rigorous standards, improving the scalability of interpretability techniques, and expanding the research scope.
@@ 9.1 Clarifying Concepts
To mature, mechanistic interpretability should embrace existing work, using established terminology rather than reinventing the wheel. Diverging terminology inhibits collaboration across disciplines. Presently, the terminology used for mechanistic interpretability partially diverges from mainstream AI research (Casper, 2023). For example, while the mainstream speaks of distributed representations (Hinton, 1984; Olah, 2023) and the goal of disentangling representations (Higgins et al., 2018; Locatello et al., 2019), the mechanistic interpretability literature refers to the same phenomenon as polysemanticity (Scherlis et al., 2023; Lecomte et al., 2023; Marshall & Kirchner, 2024) and superposition (Elhage et al., 2022b; Henighan et al., 2023). Using common language invites "accidental" contributions and prevents isolating mechanistic interpretability from broader AI research.
@@
  • Neurosymbolic Reasoning and Program Synthesis: Mechanistic interpretability aims to reverse engineer neural networks by converting their weights into human-readable algorithms. This endeavor can draw inspiration from neurosymbolic reasoning (Riegel et al., 2020) and program synthesis. Techniques like creating programs in domain-specific languages (Verma et al., 2019b; a; Trivedi et al., 2021), extracting decision trees (Zhang et al., 2019) or symbolic causal graphs (Ren et al., 2023) from neural networks align well with the goals of mechanistic interpretability. Adopting these approaches can extend the toolkit for reverse engineering AI systems.
  • Trojan Detection: Detecting deceptive models is a key motivation for inspecting model internals, as deception is not salient from observing behavior alone by definition (Casper et al., 2024). However, quantifying progress is challenging due to the lack of evidence for deception as an emergent capability in current models (Steinhardt, 2023), apart from sycophancy (Sharma et al., 2023) and theoretical evidence for deceptive inflation behavior (Lang et al., 2024). Detecting trojans or backdoors (Hubinger et al., 2024) implanted via data poisoning could serve as a helpful proxy goal and proof-of-concept. While these trojans simulate outer alignment failure (misalignment between the model’s behavior and its specified objective) rather than inner alignment failure like deceptive alignment (where an emergent sub-component optimizer within the model is misaligned with the original training objective), trojan detection still provides a practical testbed for benchmarking interpretability methods and evaluating their effectiveness quantitatively.
@@
The linear representation hypothesis treats activation directions as features (Park et al., 2023; Nanda et al., 2023b; Elhage et al., 2022b), but the emergence and necessity of linearity is unclear - is it architectural bias or inherent? Stronger theory justifying linearity’s necessity or counter-examples like autoencoders on uncorrelated data without intermediate linear layers (Elhage et al., 2022b) are needed. An alternative lens views features as polytopes from piecewise linear activations (Black et al., 2022), questioning if direction simplification suffices or added polytope complexity aids interpretability.
Polysemantic neurons are attributed to superposition compressing many features into limited neurons (Elhage et al., 2022b), but incidental redundancy without compression also causes polysemanticity (Lecomte et al., 2023; Marshall & Kirchner, 2024; McGrath et al., 2023). Understanding superposition’s role could inform mitigating polysemanticity via regularization (Lecomte et al., 2023). Superposition also raises open questions like operationalizing computation in superposition (Vaintrob et al., 2024), attention head superposition (Elhage et al., 2022b; Jermyn et al., 2023; Lieberum et al., 2023; Gould et al., 2023), representing feature clusters (Elhage et al., 2022b), connections to adversarial robustness (Elhage et al., 2022b), anti-correlated feature organization (Elhage et al., 2022b), and architectural effects (Nanda, 2023a).
@@ 9.2 Setting Standards
As the mechanistic interpretability community expands, it is essential to maintain the norm of not advancing AI capabilities while simultaneously establishing metrics necessary for the field’s progress (Räuker et al., 2023). Researchers should prioritize developing comprehensive tools for analyzing the worst-case performance of AI systems, ensuring robustness and reliability in critical applications. This includes focusing on adversarial tasks, such as backdoor detection and removal (Lamparth & Reuel, 2023; Hubinger et al., 2024; Wu et al., 2022), and evaluating the accuracy of explanations in producing adversarial examples (Goldowsky-Dill et al., 2023).
@@
To objectively evaluate interpretability methods, developing well-defined metrics and standardized benchmarks assessing aspects like feature attribution accuracy, circuit explanation comprehensiveness, and practical utility across diverse models/tasks is crucial (Räuker et al., 2023). Algorithmic testbeds are also essential for evaluating faithfulness (Jacovi & Goldberg, 2020) and falsifiability (Leavitt & Morcos, 2020) of techniques. Tools like Tracr (Lindner et al., 2023) can provide ground truth labels for benchmarking search methods (Goldowsky-Dill et al., 2023). Toy models studying superposition in computation (Vaintrob et al., 2024), transformers on algorithmic tasks can quantify sparsity and test intrinsic methods. Replacing components with hypothesized circuits (Quirke et al., 2024) should be the goal for comprehensive evaluation.
9.3 Scaling Up
@@
9.4 Expanding the Scope
Interpretability Across Training
While mechanistic interpretability of final trained models is a prerequisite, the field should also advance interpretability before and during training by studying learning dynamics (Nanda, 2022b; Elhage et al., 2022b; Hubinger, 2022). This includes tracking neuron development (Liu et al., 2021), analyzing neuron set changes with scale (Michaud et al., 2023), and investigating emergent computations (Quirke & Barez, 2023). Studying phase transitions could yield safety insights like reward hacking risks (Olsson et al., 2022).
Multi-Level Analysis
@@ 9.4 Expanding the Scope
New Frontiers: Vision, Multimodal, and Reinforcement Learning Models
@@
Reinforcement learning (RL) is also a crucial frontier given its role in advanced AI training via techniques like reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022), despite potentially posing significant safety risks (Bereska & Gavves, 2023; Casper et al., 2023a). Interpretability of RL should investigate reward/goal representations (TurnTrout et al., 2023; Colognese & Jozdien, 2023; Colognese, 2023; Bloom & Colognese, 2023), study circuitry changes from alignment algorithms (Jain et al., 2023; Lee et al., 2024), and explore emergent subgoals or proxies (Hubinger et al., 2019; Ivanitskiy et al., 2023).

### v2

@@ 8 Future Directions
Given the current limitations and challenges, several key research problems emerge as critical for advancing mechanistic interpretability. These problems span four main areas: emphasizing conceptual clarity Section 8.1, establishing rigorous standards Section 8.2, improving the scalability of interpretability techniques Section 8.3, and expanding the research scope Section 8.4. Each subsection presents specific research questions and challenges that need to be addressed to move the field forward.
@@ 8.1 Clarifying Concepts
To mature, mechanistic interpretability should embrace existing work, using established terminology rather than reinventing the wheel. Diverging terminology inhibits collaboration across disciplines. Presently, the terminology used for mechanistic interpretability partially diverges from mainstream AI research (Casper, 2023). For example, while the mainstream speaks of distributed representations (Hinton, 1984; Olah, 2023) and the goal of disentangled representations (Higgins et al., 2018; Locatello et al., 2019), the mechanistic interpretability literature refers to the same phenomenon as polysemanticity (Scherlis et al., 2023; Lecomte et al., 2023; Marshall & Kirchner, 2024) and superposition (Elhage et al., 2022b; Henighan et al., 2023). Using common language invites "accidental" contributions and prevents isolating mechanistic interpretability from broader AI research.
@@
  • Neurosymbolic Reasoning and Program Synthesis: Mechanistic interpretability aims for reverse engineering neural networks by converting their weights into human-readable algorithms. This endeavor can draw inspiration from neurosymbolic reasoning (Riegel et al., 2020) and program synthesis. Techniques like creating programs in domain-specific languages (Verma et al., 2019b; a; Trivedi et al., 2021), extracting decision trees (Zhang et al., 2019) or symbolic causal graphs (Ren et al., 2023) from neural networks align well with the goals of mechanistic interpretability. Adopting these approaches can extend the toolkit for reverse engineering AI systems.
  • Trojan Detection: Detecting deceptive alignment models is a key motivation for inspecting model internals, as – by definition – deception is not salient from observing behavior alone (Casper et al., 2024). However, quantifying progress is challenging due to the lack of evidence for deception as an emergent capability in current models (Steinhardt, 2023), apart from sycophancy (Sharma et al., 2023) and theoretical evidence for deceptive inflation behavior (Lang et al., 2024). Detecting trojans (or backdoors) (Hubinger et al., 2024) implanted via data poisoning could be a proxy goal and proof-of-concept. These trojans simulate outer misalignment (where the model’s behavior is misaligned with the specified reward function or objectives due to poorly defined or incorrect reward signals) rather than inner misalignment such as deceptive alignment (where the model appears aligned with the specified objectives but internally pursues different, misaligned goals). Moreover, activating a trojan typically results in an immediate change of behavior, while deception can be subtle, gradual, and, at first, entirely internal. Nevertheless, trojan detection can still provide a practical testbed for benchmarking interpretability methods (Maloyan et al., 2024).
@@
The linear representation hypothesis treats activation directions as features (Park et al., 2023a; Nanda et al., 2023b; Elhage et al., 2022b), but the emergence and necessity of linearity is unclear – is it architectural bias or inherent? Stronger theory justifying linearity’s necessity or counter-examples like autoencoders on uncorrelated data without intermediate linear layers (Elhage et al., 2022b) are needed. An alternative lens views features as polytopes from piecewise linear activations (Black et al., 2022), questioning if direction simplification suffices or added polytope complexity aids interpretability.
The superposition hypothesis suggests that polysemantic neurons arise from the network compressing and representing many features within its limited set of neurons (Elhage et al., 2022b), but polysemanticity can also occur incidentally due to redundancy (Lecomte et al., 2023; Marshall & Kirchner, 2024; McGrath et al., 2023). Understanding superposition’s role could inform mitigating polysemanticity via regularization (Lecomte et al., 2023). Superposition also raises open questions like operationalizing computation in superposition (Vaintrob et al., 2024), attention head superposition (Elhage et al., 2022b; Jermyn et al., 2023; Lieberum et al., 2023; Gould et al., 2023), representing feature clusters (Elhage et al., 2022b), connections to adversarial robustness (Elhage et al., 2022b), anti-correlated feature organization (Elhage et al., 2022b), and architectural effects (Nanda, 2023a).
@@ 8.2 Setting Standards
As the mechanistic interpretability community expands, it is essential to maintain the norm of not advancing AI capabilities while simultaneously establishing metrics necessary for the field’s progress (Räuker et al., 2023). Researchers should prioritize developing comprehensive tools for analyzing the worst-case performance of AI systems, ensuring robustness and reliability in critical applications. This includes focusing on adversarial tasks, such as backdoor detection and removal (Lamparth & Reuel, 2023; Hubinger et al., 2024; Wu et al., 2022a), and evaluating the accuracy of explanations in producing adversarial examples (Goldowsky-Dill et al., 2023).
@@
A central challenge in mechanistic interpretability is the lack of rigorous evaluation methods. Relying solely on intuition can lead to conflating hypotheses with conclusions, resulting in cherry-picking and optimizing for best-case rather than average or worst-case performance (Rudin, 2019; Miller, 2019; Räuker et al., 2023; Casper, 2023). Current ad hoc practices and proxy measures (Doshi-Velez & Kim, 2017) risk over-optimization (Goodhart’s law – When a measure becomes a target, it ceases to be a good measure). Distinguishing correlation from causation is crucial, as interpretability illusions demonstrate that visualizations may be meaningless without causal linking (Bolukbasi et al., 2021; Friedman et al., 2023a; Olah et al., 2017).
To advance the field, rigorous evaluation methods are needed. These should include: (i) assessing out-of-distribution inputs, as most current methods are only valid for specific examples or datasets (Räuker et al., 2023; Ilyas et al., 2019; Mu & Andreas, 2020; Casper et al., 2023c; Burns et al., 2023); (ii) controlling systems through edits, such as implanting or removing trojans (Mazeika et al., 2022) or targeted editing (Ghorbani & Zou, 2020; Dai et al., 2022; Meng et al., 2022a; b; Bau et al., 2018; Hase et al., 2023); and (iii) replacing components with simpler reverse-engineered alternatives (Lindner et al., 2023). Algorithmic testbeds are essential for evaluating faithfulness (Jacovi & Goldberg, 2020) and falsifiability (Leavitt & Morcos, 2020). Tools like Tracr can provide ground truth labels for benchmarking search methods (Goldowsky-Dill et al., 2023), while toy models studying superposition in computation (Vaintrob et al., 2024) and transformers on algorithmic tasks can quantify sparsity and test intrinsic methods. The ultimate goal should be comprehensive evaluation through replacing components with hypothesized circuits (Quirke et al., 2024).
8.3 Scaling Techniques
@@
8.4 Expanding Scope
Interpretability Across Training.
While mechanistic interpretability of final trained models is a prerequisite, the field should also advance interpretability before and during training by studying learning dynamics (Nanda, 2022b; Elhage et al., 2022b; Hubinger, 2022). This includes tracking neuron development (Liu et al., 2021), analyzing neuron set changes with scale (Michaud et al., 2023), and investigating emergent computations (Quirke & Barez, 2023). Studying phase transitions could yield safety insights for reward hacking risks (Olsson et al., 2022).
Multi-Level Analysis.
@@ 8.4 Expanding Scope
New Frontiers: Vision, Multimodal, and Reinforcement Learning Models.
@@
Reinforcement learning (RL) is also a crucial frontier given its role in advanced AI training via techniques like reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022), despite potentially posing significant safety risks (Bereska & Gavves, 2023; Casper et al., 2023a). Interpretability of RL should investigate reward/goal representations (TurnTrout et al., 2023; Colognese & Jozdien, 2023; Colognese, 2023; Bloom & Colognese, 2023), study circuitry changes from alignment algorithms (Prakash et al., 2024; Jain et al., 2023; Lee et al., 2024), and explore emergent subgoals or proxies (Hubinger et al., 2019; Ivanitskiy et al., 2023) such as internal reward models (Marks et al., 2023b).
@@ Acknowledgements
Glossary
circuits
Sub-graphs within neural networks consisting of and the weights connecting them. Circuits can be thought of as computational primitives that perform understandable operations to produce (ideally interpretable) features from prior (ideally interpretable) features. Examples include circuits for detecting curves at specific orientations (Cammarata et al., 2020; 2021), continuing repeated patterns in text (Olsson et al., 2022), and resolving anaphoric references (Wang et al., 2023). While circuits can involve clearly interpretable features, the definition allows for intermediate representations that are less easily interpretable.
concepts
An abstract idea or representation derived from observations of the world. Concepts refer to the natural abstractions that a cognitive system, like a neural network, aims to capture and represent through its learned features, which may or may not align perfectly with human-defined concepts.
deceptive alignment
When a misaligned model aims to appear aligned to gain more power to take control once sufficiently powerful.
deceptive inflation
Theoretical result on deceptive behavior: policies produce trajectories that look better than they actually are from the human’s perspective with limited observations to get higher reward signals during training. This deceptive behavior arises in reinforcement learning from human feedback when the human provides feedback based only on partial observations of the trajectories, while the policy has full state information during training (Lang et al., 2024).
disentangled
In disentangled representations, individual dimensions or components correspond to distinct, independent factors of variation in the data, rather than representing a tangled mixture of these factors.
eliciting latent knowledge
Developing strategies to make a machine learning model explicitly report latent facts or knowledge embedded in its parameters, especially in cases where the model’s output is untrusted (Christiano et al., 2021). This involves finding patterns in neural network activations that track the true state of the world (Mallen & Belrose, 2023).
features
The fundamental units of how neural networks encode knowledge, which cannot be further decomposed into smaller, distinct . Features are core components of a neural network’s representation, analogous to how cells form the fundamental unit of biological organisms (Olah et al., 2020). The hypothesis suggests an alternative definition: that features correspond to the concepts that a larger, sparser network with sufficient capacity would learn to represent with individual () neurons (Olah et al., 2020; Bricken et al., 2023).
grokking
"Grokking refers to the surprising phenomenon of delayed generalization where neural networks, on certain learning problems, generalize long after overfitting their training set." (Liu et al., 2022a)
hydra effect
The phenomenon where models can internally self-repair and maintain capabilities even when key components are ablated, making it challenging to identify the relevant components underlying a particular behavior (McGrath et al., 2023).
inner misalignment
Inner misalignment, or goal misgeneralization, occurs when an AI system develops goals or behaviors during training that are misaligned with the intended objectives despite a correctly specified reward signal.
internal world models
Internal causal environment models formed within neural networks, implicitly emerging as a by-product of prediction (e.g., in large language models).
irreducible
We adopt the notion of features as the fundamental units of neural network representations, such that features cannot be further decomposed into smaller, distinct factors. To make this more precise, we can formalize the definition of features as irreducible input patterns following Engels et al. (2024): A feature $f$ of sparsity $s$ is a function that maps a subset of the input space (with probability $01-s>$ 0) into a higher-dimensional representational space. We say the feature is active on this subset. A feature $f$ is reducible into features $a$ and $b$ if there exists a transformation that decomposes $f$ into $a$ and $b$, such that the transformed distribution $p(a,b)$ is either:
  1. 1.
    Separable: $p(a,b)=p(a)p(b)$
  2. 2.
    A mixture: $p(a,b)=wp_{1}(a,b)+(1-w)p_{2}(a,b)$ where $p_{1}$ is lower-dimensional.
Features are defined as irreducible patterns that cannot be decomposed into separable or mixture distributions via such transformations. This formalizes the notion that features form the fundamental atomic units underlying neural representations. Features that can be disentangled into statistically independent components (separable) or simpler lower-dimensional factors (mixtures) are not considered the core representational primitives. The key properties are that 1) features map from the input space to higher-dimensional representational spaces, 2) features are sparse and only activated on subsets of the input, and crucially, 3) features are irreducible and cannot be expressed as transformations of other statistically independent components.
iterative distillation and amplification
A technique for training AI systems by repeatedly distilling knowledge from a larger model into a smaller one while amplifying the smaller model’s capabilities through feedback and interaction with humans.
linear representation
Features are directions in activation space, i.e., linear combinations of neurons.
machine unlearning
Techniques for removing private data or dangerous knowledge from models.
mesa-optimization
The emergence of unintended subagents within a model with their own objectives, potentially misaligned with the original training objective.
microscope AI
Systems that extract and utilize knowledge from a model without allowing the model to take autonomous actions. This involves reverse engineering a trained model to understand its learned knowledge about the world, aiming to leverage this understanding directly without deploying the model in an operational capacity.
modularity
The property of an AI system being composed of distinct, semi-independent components or submodules that can be separately understood, modified, and recombined, rather than a monolithic, opaque structure.
monosemantic
A neuron corresponding to a single concept. The intuition is that analyzing what inputs activate a given neuron reveals its associated semantic meaning or concept. In contrast to .
motifs
Repeating patterns that emerge across models and tasks, manifesting as circuits, features, or higher-level behaviors from component interactions. Examples include curve detectors, induction circuits, and branch specialization. Motifs reveal common structures and mechanisms underlying neural network intelligence.
natural abstractions
High-level summaries or descriptions of a system or environment learned and used by many cognitive systems. According to the natural abstraction hypothesis (Chan et al., 2023), a set of "natural" abstractions exist that represent redundantly encoded information in the world and tend to be learned by intelligent systems produced through local selection pressures. These natural abstractions form a relatively small, discrete set of concepts like "tree," "velocity," etc., that allow compact descriptions of the world while discarding many irrelevant low-level details.
outer misalignment
Outer misalignment, or reward hacking, occurs when the specified reward function or utility function fails to capture the desired objectives correctly. This leads the AI to optimize for behaviors that achieve high reward scores but are misaligned with the intended outcomes.
oversight
(Scalable) oversight refers to the challenge of providing reliable supervision—through labels, reward signals, or critiques—to AI models, ensuring effectiveness even as models surpass human-level performance.
polysemantic
Neurons that are associated with multiple, unrelated concepts, contradicting the interpretation of neurons as representational primitives and making it challenging to understand the information processing of neural networks. This term is derived from linguistic concepts of polysemy (Falkum & Vicente, 2015), and in the context of neural networks first introduced by Arora et al. (2018), who suggested that word embeddings of polysemous words may be stored as a of vectors representing distinct meanings. Olah et al. (2020) first used the term polysemanticity, elaborating on the concept of polysemantic neurons as a challenge for mechanistic interpretability.
prediction orthogonality
A model whose objective is prediction can simulate agents who optimize toward any objectives with any degree of optimality (janus, 2022).
privileged basis
In certain neural network representations, the basis directions formed by the individual neurons are architecturally distinguished from arbitrary directions in the activation space. This privileged basis makes it meaningful to analyze the properties and roles of individual neurons, as the architecture encourages features to align with these basis directions. Hence, a privileged basis is necessary but not sufficient for the formation of monosemantic neurons. (Elhage et al., 2022b).
representation engineering
A top-down approach to transparency research that treats representations as the fundamental unit of analysis, aiming to understand and control representations of high-level cognitive phenomena in neural networks like large language models. Representation engineering has two main areas: 1) Reading representations to probe and interpret their contents, and 2) Controlling representations to manipulate high-level concepts like honesty or morality (Zou et al., 2023).
reverse engineering
The process of deconstructing a neural network’s computations to fully understand and specify its operations. This involves breaking down the network’s functionality into explicit, interpretable components, potentially as clear and detailed as pseudocode.
reward hacking
See outer misalignment.
simulacra
The text outputs generated by a predictive model simulating the causal processes underlying text creation. These outputs simulate coherent and contextually relevant language, sometimes exhibiting agentic behaviors or goals despite the predictive model itself lacking genuine agency or intentionality. Simulacra can be either agentic, mimicking intentional and persuasive language use, or non-agentic, merely generating descriptive text without simulated goals or agency (janus, 2022; Bereska & Gavves, 2023).
simulation
The simulation hypothesis says that when scaled up sufficiently, predictive models will learn to simulate the real-world causal processes that generated their training data (janus, 2022). When these models are optimized for predictive accuracy on broad data distributions like natural language, they are incentivized to discover the underlying rules, physics, and semantics that govern the data to model and predict future observations effectively. This allows the models to go beyond just memorizing or pattern-matching their training sets, instead learning to simulate hypothetical scenarios, reason about counterfactuals, and exhibit behaviors characteristic of general intelligence – all as a byproduct of the drive for efficient compression and accurate prediction. The simulation hypothesis suggests these models will develop rich internal world models capturing the causal dynamics of the training distribution.
streetlight interpretability
Examining AI systems under only ideal conditions of maximal interpretability, risking missing critical phenomena that only emerge in more realistic and diverse contexts.
superposition
The superposition hypothesis suggests that neural networks can leverage high-dimensional spaces to represent more features than the actual count of neurons by encoding features in almost orthogonal directions (Elhage et al., 2022b).
sycophancy
The tendency of models to generate responses that align with user beliefs rather than providing truthful information. This behavior, encouraged by human feedback used in fine-tuning, is observed in state-of-the-art AI assistants across various tasks (Sharma et al., 2023). Sycophancy arises because human preference judgments often favor responses that match users’ views, leading to a preference for convincingly written sycophantic responses over correct ones.
universality
The universality hypothesis proposes the emergence of common circuits across neural network models trained on similar tasks and data distributions. A stronger form posits that these common circuits represent a set of fundamental computational motifs that neural networks gravitate towards when learning. The weaker version suggests that for a given task, dataset, and model architecture, an optimal way to solve the problem may exist, which different models will tend to converge towards, resulting in analogous circuits. The universality hypothesis implies that rather than each model learning arbitrary, unstructured representations, there is an underlying universality to the circuits that emerge, shaped by the learning task and inductive biases.
well-founded AI
Developing AI systems with provable safety guarantees about their behavior and alignment with human values through rigorous mathematical modeling and verification. (Tegmark & Omohundro, 2023; Dalrymple et al., 2024).